OpenStudy (anonymous):

Look at the pattern shown below. What number is missing from the pattern? 20, 22, 26, 34, __,82 42 52 50 Not Here 38

OpenStudy (anonymous):

the sequence is 2,4,8,16,32 which is 20,22,26,34,50,82 so: the answer is 50

OpenStudy (hoblos):

you are adding a power of 2 each time 20+2=22 22+4=26 26+8=34 34+16 = 50

OpenStudy (anonymous):

thankkkkkk you both so much!!!